# The Percept-V Challenge: Can Multimodal LLMs Crack Simple Perception Problems?

### v2

@@ Abstract
Cognitive science research treats visual perception, the ability to understand and make sense of a visual input, as one of the early developmental signs of intelligence. Its TVPS-4 framework categorizes and tests human perception into seven skills such as visual discrimination, and form constancy. Do Multimodal Large Language Models (MLLMs) match up to humans in basic perception? Even though there are many benchmarks that evaluate MLLMs on advanced reasoning and knowledge skills, there is limited research that focuses evaluation on simple perception. In response, we introduce Percept-V, a dataset containing 6000 program-generated uncontaminated images divided into 30 domains, where each domain tests one or more TVPS-4 skills. Our focus is on perception, so we make our domains quite simple and the reasoning and knowledge required for solving them are minimal. Since modern-day MLLMs can solve much more complex tasks, our a-priori expectation is that they will solve these domains very easily. Contrary to our belief, our experiments show a weak performance of SoTA proprietary and open-source MLLMs compared to very high human performance on Percept-V. We find that as number of objects in the image increases, performance goes down rather fast. Our experiments also identify the perception skills that are considerably harder for all models.
@@
Samrajnee Ghosh  and Naman Agarwal  and Hemanshu Garg  and Chinmay Mittal and Mausam and Parag Singla
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated high performance on complex tasks, leading some researchers to suggest that these models are approaching human-level intelligence (Zhou et al., 2023; King, 2023) or even achieving MENSA-level capabilities (Schregel, 2020). However, current visual benchmarks typically assess image understanding alongside reasoning and specialized domain knowledge in fields such as engineering, medicine, and art (Nguyen, 2023; Yue et al., 2024). This integrated approach makes it challenging to isolate MLLMs’ pure image understanding capabilities from their reasoning abilities and domain-specific knowledge. To our knowledge, there is little work that focuses specifically on visual perception, the fundamental ability to understand and interpret visual images.
@@
Visual perception represents a foundational cognitive ability that develops early in human cognition, preceding the acquisition of complex reasoning and knowledge-based skills (Rabindran and Madanagopal, 2020). Recent research has raised concerns about MLLMs’ perceptual capabilities (Zhang et al., 2024b; Fu et al., 2024), highlighting the need for a dedicated perception benchmark. To address this gap, we introduce Percept-V, a perception-focused dataset comprising 6000 program-generated, uncontaminated image-based problems from 30 domains, which primarily evaluate visual understanding while requiring minimal problem-solving skills and no specialized domain knowledge. Sample questions are presented in Figure 1.
@@
Visual Closure - These are two images of some sequence of shapes separated by black horizontal lines - The first image consists of a sequence of shapes and the second image consists of a sequence of outlines of shapes - Determine the number of shapes in the first image that match the outlines in the corresponding place in the second image Refer to caption correct ans: 5 GPT-4o ans: “ANSWER: 3” Visual Sequential Memory - This image contains some concentric circles in different colors. - Write down the sequence of colors seen from inside to outside - The colors belong too the following list - [black,gray,brown,maroon,red,coral,tan,orange, ivory,goldenrod,yellow,green,olive,turquoise, skyblue,lavender, purple, pink, fuschsia] - No color is repeated. Refer to caption correct ans: ivory,coral,pink,green,orange,black,fucshia, yellow,skyblue,lavender GPT-4o ans: ivory,coral,pink,green,orange,black,pink, yellow, fucshia,skyblue,lavender Visual Discrimination - These are two images containing shapes such as circles, triangles and squares - The images are identical except for the fact that the second image has a few objects missing - Count the number of missing objects in the secnd image Refer to caption correct ans: 10 GPT-4o ans: “COUNT: 5” Visual Figure Ground - This is an image with some multi-coloured shapes drawn on top of it - Identify the shapes drawn over the background image - Form a set of all the unique shapes seen in the image Refer to caption correct ans: circle,pentagon,rectangle,triangle GPT-4o ans: “circle,triangle” Visual Memory - This is an image with some shapes drawn and numbers written on top of the shapes - Determine the list of numbers written on top of the circles. Refer to caption correct ans: 1,3,6,7,9,10,14,15,16,20 GPT-4o ans: “CIRCLES:1,3,7,10,14,15” Visual Spatial Relationship - There will be an image containing a rectangle and some circles above and below it - Count the number of circles above the rectangle and below the rectangle Refer to caption correct ans: “ABOVE:10,BELOW:10” GPT-4o ans: “ABOVE:8,BELOW:10” Visual Form Constancy - This is an image containing some black circles - Count the total number of circles in the image Refer to caption correct ans: 10 GPT-4o ans: “CIRCLES:10”
@@
Figure 1: Sample questions from the Percept-V dataset illustrating tasks related to different visual perception skills as defined by the TVPS-4 framework. Each example shows the image prompt, the correct answer, and a typical incorrect response from an MLLM.
@@
Our dataset construction draws from cognitive science literature, specifically utilizing the skill classification framework from TVPS-4 (Test of Visual Perceptual Skills) (Martin and Gardner, 2006), which is widely used for human perception assessment. TVPS-4 organizes visual perceptual abilities into seven distinct categories: visual discrimination, visual memory, visual sequential memory, visual-spatial relationships, figure-ground perception, form constancy, and visual closure (detailed in Section 3).
@@
Percept-V comprises 30 domains, each designed to evaluate a combination of 1-3 TVPS-4 skill categories. Every domain features a single standardized question prompt that describes the specific task requirements. To systematically assess performance across varying levels of difficulty, each domain contains 200 images with varying problem size, which is operationalised through parameters such as the number of objects present, grid dimensions, or the number of processing steps required to complete the task. This structured approach enables us to evaluate MLLM performance as a function of visual complexity and provides insight into how they handle increasing perceptual demands.
We experiment on three state-of-the-art proprietary MLLMs – GPT-4o, O4-mini and Gemini 2.5 Flash and two open source MLLMs – Qwen 2.5 VL Instruct and DeepSeek VL2 Tiny, which include both language and reasoning models, and also perform a human study. Our findings reveal substantial gaps between MLLM and human performance on these simple visual perception tasks. Across all evaluated models, we observe consistent performance degradation as problem size increases, suggesting systematic limitations in handling visual complexity. While some models demonstrate relative strength in specific skill categories, overall performance on a skill remains remarkably similar across models, indicating broad deficits in visual understanding. Overall, our work is an important step towards understanding the visual perception limitations that must be addressed in the future to develop more robust MLLMs.
@@ 2 Related Work
Most exist benchmarks for evaluating the performance of MLLMs necessitate a combination of knowledge, reasoning and perception for good performance. The NTSEBench dataset (Pandya et al., 2024) comprises questions from the commonsense reasoning test from India. The MathVista dataset (Lu et al., 2023) examines different SoTA models on fine-grained, deep visual understanding and compositional reasoning. The MathVision benchmark (Wang et al., 2024) tests MLLMs on questions with a visual context with 16 distinct mathematical disciplines and 5 levels of difficulty.
OlympiadBench(He et al., 2024) collects advanced reasoning questions from the fields of mathematics, physics and chemistry to assess the complex reasoning abilities of MLLMs. Works like MMMU (Yue et al., 2024) evaluate MLLMs on many disciplines such as engineering, medicine, humanities, science, business and art. An extension of MMMU in Chinese called CMMMU (Zhang et al., 2024a) additionally tests the language understanding skills of MLLMs. SciBench (Wang et al., 2023) contains college-level science questions. Some datasets require interpretation of abstract figures, geometry shapes and scientific plots of ArXiv papers (Li et al., 2024), and experiments suggest that MLLMs could likely be used as scientific assistants in the future. EMMA Hao et al. (2025) is a recent work requiring cross-modal reasoning for answering problems in domains such as mathematics, physics, chemistry and coding. All these datasets, along with perception, require complex reasoning skills to answer questions, and several datasets additionally require specialized knowledge of a subject.
Some recent benchmarks specifically try to carefully disentangle the reasoning and perceptual abilities of LLMs. (Chung et al., 2025) propose MatHLENS which separately tests the perceptual, reasoning and integrating capabilites of MLLMs. In recent times, additional studies and benchmarks have been proposed that specifically highlight the lack of MLLMs’ perceptual abilities, such as Lee et al. (2025). Zhang et al. (2024b) shows that MLLMs perform poorly in small object and low quality object recognition. Fu et al. (2024) develops a benchmark of simple perceptual questions that can be hard for MLLMs. Most of these works use complex image scenes for evaluation and do not study the impact of problem size on perception tasks.
@@
Most related to our work is a very recent unpublished paper (Kanade and Ganu, 2025), which also develops a dataset in the same spirit as ours. While it is contemporaneous research, our dataset is much larger – 30 domains, compared to 12, and allows for a more fine-grained variation along problem size, making our evaluation more extensive. We ground each domain into the specific skills from TVPS-4 framework covering all 7 skills, whereas they work with only a subset of them. 43% of our domains require multiple TVPS-4 skills, which allows us to assess MLLM’s ability when applying skills in concert. Their work does not explicitly address the issue of handling multiple skills in a single domain.
@@ 3 Dataset
We now describe the details of our dataset referred to as Percept-V. Our dataset is composed of 30 different domains where each domain tests one or more basic perception skills from the TVPS-4 cognitive framework. The skills listed in the TVPS-4 framework are as follows:
@@
    Visual Spatial Relationships: - Understanding the relationships of objects within the environment.
@@
    Visual Form Constancy: The ability to know that a form or shape is the same, even if it has been made smaller/larger or has been turned around.
@@
    Visual Closure: The ability to recognize a form or object when part of the picture is missing.
@@
Each domain in our dataset consists of multiple problem instances which require a subset of TVPS-4 skills Martin and Gardner (2006) for them to be solved correctly. Table 1 lists the domains for each skill, which require the knowledge of that skill in order for its problem instances to be solved. In total, 17 domains test one skill, 8 domains have two skills, and 5 domains test 3 skills in concert.
@@
Each domain is associated with a unique question prompt. Further, each problem instance in a domain is composed of a pair $(Q,I)$, where $Q$ represents the question prompt associated with the domain, and $I$ is an image generated automatically through Python programmes that used libraries like Matplotlib, Pillow and OpenCV.111image generation for each domain is controlled by a set of hyperparameters described in A. In other words, we apply the same generic question prompt to different images resulting in different problem instances. The problem instances are categorized across multiple sizes, where size of a problem instance captures its inherent complexity, such as number of objects or number of distractors, or number of sequence of steps required to solve the problem instance. For example, in the counting_circlescounting\_circles domain shown under Visual Form Constancy in 1, the problem size is controlled by controlling the number of circles to be counted. On the other hand, in the numbered_shapesnumbered\_shapes domain under Visual Memory in 1, with increasing problem size, the number of circles as well as distractors i.e., triangles and squares increase. Further, each question prompt, $Q$ is composed of three types of prompts $(in,r,op)$, where $in$ provided the description of the input image $I$, $r$ provided details of the task to be performed and $op$ specified the answer format of the model output.
Each domain has its problem instances divided across the sizes varying from 1 to 20, with 10 instances generated for each size in each domain in the dataset. This results in 200 problem instances for each domain and a total of 6000 problem instances divided across 30 domains. Examples of problem instances from our dataset are provided in Figure 1.
@@
There are four types of question prompts in our dataset based on the type of the answer they expect. Details are listed in Table 2. The answer can be a single value (Boolean, numeric), a fixed length answer (a tuple or a list of a fixed length for all problem sizes), an ordered list (a list that has to maintain a specific sequence with its length dependent on the problem size) or an unordered set (a set with only unique values). For each answer type, the question prompt also lists a specific format in which the answer is to be outputted, so that it can be automatically evaluated. All MLLMs mostly honor the format consistently – the errors are from incorrect perception and not format mismatches.
We note all the images in our dataset are generated using automated programs. This has multiple advantages: (1) it helps control the complexity of the problem instances created, and (2) it helps avoid any potential contamination issues, which has been pointed out as a serious concern in some of the existing benchmarks (Shojaee et al., 2025). Further details of the images in each domain along with specific examples, as well as our automated generative process, is described in the Appendix.
@@ 4 Experiments
Through our experiments, we try to answer the following questions:
  1. 1.
    How well do the SOTA open as well as proprietary LLMs perform on Percept-V?
  2. 2.
    Do some LLMs perform better than others especially for a subset of skills?
  3. 3.
    How does the LLM performance vary across variation in skills: are some skills harder than others?
  4. 4.
    How does the LLM performance vary with increasing size of the problem instances in different skills?
  5. 5.
    Finally, how does the LLM performance compare with human performance on Percept-V?
In order to answer these questions, we perform our experimental comparison using three different state-of-the-art proprietary LLMs: (1) GPT-4o (2) Gemini 2.5 Flash (3) o4-mini, and two open source MLLMs: (1) Qwen-2.5-7B-Instruct (2) Deepseek-VL2-Tiny. We note that among these, o4-mini and Gemini 2.5 Flash are reasoning/thinking models. For open source models, all our experiments were run on an NVidia A100 GPU. More details about model parameters and GPU runtime are presented in A. The hyperparameters like temperature and top_p are both kept as 0.0 to produce more deterministic reproducible answers from models like Gemini, GPT-4o and Qwen (for the remaining models, the temperature could not be made 0.0 due to model restrictions). We run all experiments in a zero-shot mode to keep the costs under control; further, our preliminary experiments on a subset of models suggest that few shot prompting does not yield any significant gains for our domains.
For each problem size of each domain, all 10 candidate problems are evaluated on all models and the average performance on these 10 problems is considered as the model accuracy for that problem size. For each LLM for a given domain, we use exactly the same prompt. This prompt is constructed by appending the question prompt followed by details of output format. Detailed examples of prompts are given in the Appendix.
Skill / Domain GPT-4o o4-mini Gemini Qwen DeepSeek Avg. Acc.
@@
Visual Discrimination 23.70 55.30 51.25 11.50 4.25 29.20
Visual Memory 33.55 47.55 46.55 19.59 1.50 29.75
Visual Sequential Memory 21.50 47.50 41.92 4.67 5.42 24.20
Visual Figure Ground 25.42 63.25 62.83 7.58 0.08 31.83
Visual Form Constancy 27.30 48.90 50.30 25.60 2.50 30.92
Visual Closure 18.40 40.20 34.60 4.60 6.50 20.86
Visual Spatial Relationship 50.36 62.64 69.79 27.79 25.21 47.16
Average of skills 28.60 52.19 51.03 14.48 6.49 37.04
@@
change colour 18.50 32.00 23.50 8.00 4.50 17.30
circle boxes 33.00 34.00 28.00 12.50 10.50 23.60
circle location 41.50 46.00 57.50 14.50 6.00 33.10
circle right triangle 82.00 99.50 99.00 25.00 50.00 71.10
colours present 1.50 4.50 4.50 0.00 0.00 2.10
comparing size 36.50 31.00 55.00 5.00 0.00 25.50
count coloured circles 44.00 37.50 44.50 39.00 5.00 34.00
counting circles 57.50 55.00 55.50 42.50 4.00 42.90
counting locations 20.50 36.50 33.50 15.50 0.00 21.20
counting shapes 44.00 58.50 61.00 41.00 0.00 40.90
cross and knots 57.50 100.00 96.00 23.50 47.50 64.90
graph counting 31.00 33.50 36.00 26.00 0.00 25.30
grid path 0.50 53.00 23.50 0.00 0.00 15.40
identifying shapes 69.00 92.50 47.00 41.50 0.00 50.00
inside circles 87.50 95.50 93.00 96.50 73.00 89.10
layered colours 25.00 33.50 34.00 0.50 0.00 18.60
layered shapes 11.50 20.00 18.50 9.50 0.00 11.90
list colours 18.00 35.50 58.00 1.50 0.50 22.70
list shapes 70.50 90.50 92.00 21.00 0.00 54.80
locate circles colour 11.50 100.00 82.00 6.00 0.00 39.90
locate circles shape 16.00 100.00 92.50 7.00 0.00 43.10
match outline 23.00 84.50 72.50 7.50 15.00 40.50
match shadow 32.50 63.00 48.00 5.50 17.50 33.30
maze solving 27.00 30.00 54.50 14.50 0.00 25.20
mirror image 12.50 29.00 24.00 21.50 5.00 18.40
numbered shapes 12.00 75.50 88.50 8.00 0.00 36.80
sort circles 24.00 34.00 30.00 20.00 0.50 21.70
sort lines 22.00 27.00 22.50 5.00 0.00 15.30
vanishing objects 9.00 23.00 25.50 10.00 5.00 14.50
water image 10.50 26.50 22.50 15.00 3.50 15.60
Full Percept-V dataset 31.65 52.70 50.75 18.10 8.25 32.29

### v3

@@ Abstract
Cognitive science research treats visual perception, the ability to understand and make sense of a visual input, as one of the early developmental signs of intelligence. Its TVPS-4 framework categorizes and tests human perception into seven skills such as visual discrimination, and form constancy. Do Multimodal Large Language Models (MLLMs) match up to humans in basic perception? Even though many benchmarks evaluate MLLMs on advanced reasoning and knowledge skills, there is limited research that focuses evaluation on simple perception. In response, we introduce Percept-V, a dataset containing 6000 program-generated uncontaminated images divided into 30 domains, where each domain tests one or more TVPS-4 skills. Our focus is on perception, so we make our domains quite simple and the reasoning and knowledge required for solving them are minimal. Since modern-day MLLMs can solve much more complex tasks, our a-priori expectation is that they will solve these domains very easily. Contrary to our belief, our experiments show a weak performance of SoTA proprietary and open-source MLLMs compared to very high human performance on Percept-V. We find that as number of objects in the image increases, performance goes down rather fast. Our experiments also identify the perception skills that are considerably harder for all models.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated high performance on complex tasks, leading some researchers to suggest that these models are approaching human-level intelligence (Zhou et al., 2023; King, 2023) or even achieving MENSA-level capabilities (Schregel, 2020). However, current visual benchmarks typically assess image understanding alongside reasoning and specialized domain knowledge in fields such as engineering, medicine, and art (Nguyen, 2023; Yue et al., 2024). This integrated approach makes it challenging to isolate MLLMs’ visual perception capabilities from their reasoning abilities and domain-specific knowledge.
@@
Visual perception represents a foundational cognitive ability that precedes the acquisition of complex reasoning and knowledge-based skills in humans (Rabindran and Madanagopal, 2020). Recent research has raised concerns about MLLMs’ perceptual capabilities (Zhang et al., 2024b; Fu et al., 2024), highlighting the need for a dedicated perception benchmark. To address this gap, we introduce Percept-V, a dataset comprising 6000 program-generated, uncontaminated image-based problems from 30 domains, which primarily evaluate visual understanding while requiring minimal problem-solving skills and no specialized domain knowledge. See Figure 1 for sample questions.
@@
Visual Closure Determine the number of shapes in the rows of the first column that match the outlines in the corresponding rows in the second column. Refer to caption correct ans: 5 GPT-4o ans: “ANSWER: 3” Visual Sequential Memory This image contains some concentric circles in different colors. Write down the sequence of colors seen from inside to outside. Refer to caption correct ans: orange, yellow, green, tan, purple, gray GPT-4o ans: orange, yellow, green, tan, purple, gray Visual Discrimination The images are identical except for the fact that the lower image has a few objects missing. Count the number of missing objects in the lower image. Refer to caption correct ans: 10 GPT-4o ans: “COUNT: 5” Visual Figure Ground Identify the shapes drawn over the background image. Form a set of all the unique shapes seen in the image. Refer to caption correct ans: circle,pentagon,rectangle,triangle GPT-4o ans: “circle,triangle” Visual Memory This is an image with some shapes drawn and numbers written on top of the shapes. Determine the list of numbers written on top of the circles. Refer to caption correct ans: 1,3,6,7,9,10,14,15,16,20 GPT-4o ans: “CIRCLES:1,3,7,10,14,15” Visual Spatial Relationship Count the number of circles above the rectangle and below the rectangle Refer to caption correct ans: “ABOVE:10,BELOW:10” GPT-4o ans: “ABOVE:8,BELOW:10” Visual Form Constancy Count the total number of circles in the image. Refer to caption correct ans: 10 GPT-4o ans: “CIRCLES:10”
@@
Figure 1: Sample questions from the Percept-V dataset illustrating tasks related to different visual perception skills as defined by the TVPS-4 framework. Each example shows the summarised image prompt, the correct answer, and a typical incorrect response from an MLLM.
@@
Our dataset construction draws from cognitive science literature, specifically utilizing the skill classification framework from TVPS-4 (Test of Visual Perceptual Skills) (Martin and Gardner, 2006), which is widely used for human perception assessment. TVPS-4 organizes visual perceptual abilities into seven distinct categories such as visual discrimination, visual-spatial relationships, and form constancy (detailed in Section 3).
@@
Percept-V comprises 30 domains, each designed to evaluate a combination of 1-3 TVPS-4 skill categories. Every domain features a single standardized question prompt that describes the specific task requirements. Each instance asks a perception question over an image that generally has simple objects such as circles and triangles, in different colors and sizes. The use of basic objects is similar to benchmarks in robotics Li et al. (2024a), and mathematics Wang et al. (2024).
To systematically assess performance across varying levels of difficulty, each domain contains 200 images with varying problem size, which is operationalised through the number of objects present, grid dimensions, or the number of processing steps required to complete the task. This enables us to evaluate MLLM performance as a function of visual complexity and provides insight into how they handle increasing perceptual demands.
We experiment on four state-of-the-art proprietary MLLMs – GPT-5-mini, GPT-4o, o4-mini and Gemini 2.5 Flash and two open source MLLMs – Qwen 2.5 VL Instruct and DeepSeek VL2 Tiny, which include both language and reasoning models, and also perform a human study. Our findings reveal substantial gaps between MLLM and human performance on these simple visual perception tasks. Across all evaluated models, we observe consistent performance degradation as problem size increases, suggesting systematic limitations in handling visual complexity. While some models demonstrate relative strength in specific skill categories, overall performance on a skill remains remarkably similar across models, indicating broad deficits in visual understanding. Overall, our work is an important step towards understanding the visual perception limitations that must be addressed in the future to develop more robust MLLMs.
@@ 2 Related Work
Most existing benchmarks for evaluating the performance of MLLMs necessitate a combination of knowledge, reasoning and perception for good performance. Alongside perception, the NTSEBench dataset (Pandya et al., 2024) tests commonsense reasoning, the MathVista dataset (Lu et al., 2023) evaluates compositional reasoning, and the MathVision benchmark (Wang et al., 2024) includes questions from 16 mathematical disciplines.
In a similar spirit, OlympiadBench (He et al., 2024) collects advanced reasoning questions from mathematics, physics and chemistry. MMMU, CMMMU and EMMA (Yue et al., 2024; Zhang et al., 2024a; Hao et al., 2025) evaluate MLLMs on many disciplines such as engineering, medicine, humanities, science, business and art. SciBench (Wang et al., 2023) contains college-level science questions. Some datasets require interpretation of abstract figures, geometry shapes and scientific plots of ArXiv papers (Li et al., 2024b). All these datasets, along with perception, require complex reasoning skills, and several datasets additionally require specialized knowledge of a subject.
Some recent benchmarks specifically try to disentangle reasoning and perception. Chung et al. (2025) propose MatHLENS which separately tests the perceptual, reasoning and integrating capabilites of MLLMs. In recent times, some studies specifically highlight the lack of MLLMs’ perceptual abilities, e.g., Lee et al. (2025). Zhang et al. (2024b) shows that MLLMs perform poorly in small object and low quality object recognition. Fu et al. (2024) develops a benchmark of simple perceptual questions that can be hard for MLLMs. Most of these works use complex image scenes for evaluation and do not study the impact of problem size on perception tasks.
@@
Most related to our work is a very recent unpublished paper (Kanade and Ganu, 2025), which also develops a dataset in the same spirit as ours. While it is contemporaneous research, our dataset is much larger – 30 domains, compared to 12, and allows for a more fine-grained variation along problem size, making our evaluation more extensive. We ground each domain into the specific skills from TVPS-4 framework covering all 7 skills, whereas they work with only a subset. 43% of our domains require multiple skills, which allows us to assess MLLM’s ability when applying skills in concert. Their work does not explicitly address this.
@@ 3 Dataset
We now describe the details of our dataset, which we name Percept-V. It has 30 different domains where each domain tests one or more basic perception skills from the TVPS-4 cognitive framework Martin and Gardner (2006). The skills listed in the TVPS-4 framework are as follows:
@@
    Visual Spatial Relationships: Understanding the relationships between objects.
@@
    Visual Form Constancy: The ability to know that a form or shape is the same, even if it has been made smaller/larger or has been rotated.
@@
    Visual Closure: The ability to recognize a form, when a part of the picture is missing.
@@
Each domain has multiple problem instances with varying perceptual complexity. Table 1 lists, for each skill, the domains that require that skill to solve the problem instances correctly. In total, 17 domains test one skill, 8 domains have two skills, and 5 domains test 3 skills in concert.
Skills associated with visual memory and visual sequential memory typically evaluate the ability to memorize and recall a detail once it has been presented and then removed from the field of view. We note that this kind of set-up is not possible for modern day MLLMs since anything once shown is always present in model’s memory. Therefore, specifically, for these skills, our domains can be seen as testing the skill of scanning and transcribing (which is a subset of the original skill), and our mapping to the skills of visual (sequential) memory, is approximate. Nevertheless, even in this set-up, which is presumably easier than the original skills that additionally require memory, the performance of MLLMs is quite low (Table 3), pointing a fundamental limitation in their basic perception skills.
@@
Each domain is associated with a unique question prompt. Further, each problem instance in a domain is composed of a pair $(Q,I)$, where $Q$ represents the question prompt associated with the domain, and $I$ is an image generated automatically through Python programs.111image generation for each domain is controlled by a set of hyper-parameters described in Appendix A In other words, we apply the same generic question prompt to different images resulting in different problem instances. The instances vary in sizes, where size of a problem instance captures its inherent complexity, such as number of objects or number of distractors or number of sequence of steps required to solve the instance. Each domain has its instances divided across the sizes varying from 1 to 20, with 10 instances generated for each size in each domain in the dataset. This results in 200 problem instances for each domain and a total of 6000 instances divided across 30 domains. Examples of instances from our dataset are provided in Figure 1. For example, in the counting_circlescounting\_circles domain, the problem size is controlled by controlling the number of circles to be counted. Similarly, in the numbered_shapesnumbered\_shapes domain, the number of circles, and the number of distractor objects, i.e., triangles and squares, increase with increasing problem size.
@@
Each question prompt, $Q$ is composed of three types of prompts $(in,r,op)$, where $in$ provides the description of the input image $I$, $r$ provides details of the task to be performed and $op$ specifies the answer format of the model output. There are four types of question prompts in our dataset based on the type of the answer they expect. Details are listed in Table 2. The answer can be a single value (Boolean, numeric), a fixed length answer (a tuple or a list of a fixed length for all problem sizes), an ordered list (a list that has to maintain a specific sequence with its length dependent on the problem size) or an unordered set (a set with only unique values). For each answer type, the question prompt also lists a specific format in which the answer is to be outputted, so that it can be automatically evaluated. All MLLMs mostly honor the format consistently – the errors are from incorrect perception and not format mismatches (see Section 4.5).
All the images in our dataset are generated using automated scripts. Every domain has a separate script to generate images specifically for that domain. The script takes a list of problem sizes and number of instances of each problem size to be generated as its arguments. The script also randomizes the position of objects in the image to ensure the generation of a new instance. Each domain uses only simple, basic shapes like circles, triangles, squares, and pentagons, and structures like grids, rows, and rectangular mazes, which are produced using standard Python libraries like Pillow, Matplotlib, and OpenCV. The answers for each problem instance are also computed during the data generation process, and hence, there is no scope for error or ambiguity in the gold answer. We note that our methodology of generating the dataset using automated scripts as detailed above has multiple advantages: (1) it helps control the complexity of the created problem instances and (2) it helps avoid any potential contamination issues, pointed out as a serious concern in some of the existing benchmarks (Shojaee et al., 2025).
@@ 4 Experiments
Through our experiments, we try to answer the following questions: (1) How well do the SoTA open as well as proprietary MLLMs perform on Percept-V? (2) Do some MLLMs perform better than others especially for a subset of skills? (3) How does an MLLM’s performance vary across variation in skills: are some skills harder than others? (4) How does the MLLM performance vary with increasing size of the problem instances in different skills? (5) Finally, how does the MLLM performance compare with human performance on Percept-V?
In order to answer these questions, we perform our experimental comparison using four different state-of-the-art proprietary LLMs: (1) GPT-4o (2) Gemini 2.5 Flash (3) o4-mini, (4) GPT-5-mini and two open source MLLMs: (1) Qwen-2.5-7B-Instruct (2) Deepseek-VL2-Tiny. We note that among these, GPT-5-mini, o4-mini and Gemini 2.5 Flash are reasoning/thinking models, and all, except Qwen 2.5 VL Instruct, are equipped with ’think with image’ capabilities – i.e., these models leverage visual information as intermediate steps in their thought process Su et al. (2025).
For open source models, all our experiments were run on an NVidia A100 GPU. More details about model parameters and GPU runtime are presented in Appendix A. The hyperparameters like temperature and top_p are both kept as 0.0 to produce more deterministic reproducible answers from models like GPT-5-mini, Gemini, GPT-4o and Qwen (for the remaining models, the temperature could not be made 0.0 due to model restrictions). We run all experiments primarily in a zero-shot mode to keep the costs under control. We further experiment with one-shot prompting, but that does not yield any significant gains for our domains. For each LLM for a given domain, we use exactly the same prompt. This prompt is constructed by appending the question prompt followed by details of output format. Detailed examples of prompts are given in the appendix.
Skill / Domain GPT-5-mini GPT-4o o4-mini Gemini Qwen DeepSeek Avg. Acc.
@@
Visual Discrimination 55.0 23.7 55.3 51.25 11.5 4.1 33.47
Visual Memory 49.95 33.55 47.55 46.55 19.59 1.32 33.08
Visual Sequential Memory 43.5 21.5 47.5 41.92 4.67 5.25 27.39
Visual Figure Ground 64.75 25.42 63.25 62.83 7.58 0.08 37.32
Visual Form Constancy 58.7 27.3 48.9 50.3 25.6 3.0 35.63
Visual Closure 44.88 23.0 50.25 43.25 5.75 7.88 29.16
Visual Spatial Relationship 66.64 50.36 62.64 69.79 27.79 17.64 49.14
Average of skills 54.77 29.26 53.63 52.27 14.64 5.61 35.03
@@
change_colour 32.0 18.5 32.0 23.5 8.0 4.5 19.75
circle_boxes 30.0 33.0 34.0 28.0 12.5 10.5 24.67
circle_location 56.5 41.5 46.0 57.5 14.5 5.5 36.92
circle_right_triangle 100.0 82.0 99.5 99.0 25.0 50.0 75.92
colours_present 3.5 1.5 4.5 4.5 0.0 0.0 2.33
comparing_size 27.0 36.5 31.0 55.0 5.0 0.0 25.75
count_coloured_circles 51.0 44.0 37.5 44.5 39.0 2.5 36.42
counting_circles 68.0 57.5 55.0 55.5 42.5 5.5 47.33
counting_locations 44.5 20.5 36.5 33.5 15.5 0.0 25.08
counting_shapes 72.5 44.0 58.5 61.0 41.0 0.0 46.17
cross_and_knots 99.5 57.5 100.0 96.0 23.5 0.0 62.75
graph_counting 39.0 31.0 33.5 36.0 26.0 0.0 27.58
grid_path 54.5 0.5 53.0 23.5 0.0 0.0 21.92
identifying_shapes 79.5 69.0 92.5 47.0 41.5 0.0 54.92
inside_circles 96.0 87.5 95.5 93.0 96.5 68.0 89.42
layered_colours 35.0 25.0 33.5 34.0 0.5 0.0 21.33
layered_shapes 24.0 11.5 20.0 18.5 9.5 0.0 13.92
list_colours 41.0 18.0 35.5 58.0 1.5 0.5 25.75
list_shapes 88.5 70.5 90.5 92.0 21.0 0.0 60.42
locate_circles_colour 100.0 11.5 100.0 82.0 6.0 0.0 49.92
locate_circles_shape 100.0 16.0 100.0 92.5 7.0 0.0 52.58
match_outline 66.0 23.0 84.5 72.5 7.5 17.0 45.08
match_shadow 54.5 32.5 63.0 48.0 5.5 14.5 36.33
maze_solving 43.0 27.0 30.0 54.5 14.5 0.0 28.17
mirror_image 33.5 12.5 29.0 24.0 21.5 5.0 20.92
numbered_shapes 81.5 12.0 75.5 88.5 8.0 1.0 44.42
sort_circles 36.5 24.0 34.0 30.0 20.0 0.0 24.08
sort_lines 26.5 22.0 27.0 22.5 5.0 0.0 17.17
vanishing_objects 35.0 9.0 23.0 25.5 10.0 5.0 17.92
water_image 38.0 10.5 26.5 22.5 15.0 3.5 19.33
Full Percept-V Dataset 55.22 31.65 52.7 50.75 18.1 6.43 35.81